# Online Flow Time Minimization: Tight Bounds for Non-Preemptive Algorithms

### v1

@@ 8.3 Generalization to Multiple Machines with Kill-and-Restart
In each busy gadget, if $c=0$, then schedule large job 0 at 0 and large job 11 at 33; if $c=1$, then schedule large job 11 at 0 and large job 0 at 44. In this way, no large jobs intersect with any small jobs. As a result, the total flow time in this gadget is at most 99 (large jobs) +1+1 (fixed small jobs) +1+1 (random small jobs). Therefore, $\textup{{OPT}}\leq 11mk$ by summing up all gadgets. ∎
@@
Note that before time 22, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time 22 is not related to the result of $c_{b}$.
  • Let $x_{2}$ be the number of jobs with processing time 22 started in [0.5, 2)[-0.5,\;2).
  • Let $x_{3}$ be the number of jobs with processing time 33 started in [0.5, 2)[-0.5,\;2).
After time 22, the algorithm scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
  • Left case (w.p. $0.5$): None of the $x_{3}$ jobs can finish without intersection, and at most $x_{2}^{(L)}=\min\{m,x_{2}\}$ out of the $x_{2}$ jobs can finish without intersection. Let $y_{2}^{(L)}$ be the number of jobs with processing time 22 starting in $[2,6.5)$ that finish without intersection, and $y_{3}^{(L)}$ be the number for processing time 33. At most $m$ jobs out of $y_{2}^{(L)}+y_{3}^{(L)}$ can finish without intersection.
  • Right case (w.p. $0.5$): At most $m$ out of these $x_{2}+x_{3}$ jobs could end up without intersection with small jobs; let $x_{2}^{(R)}$ and $x_{3}^{(R)}$ be the number of such jobs, respectively. At most $m$ jobs with processing time 22 released in $[2,6.5)$ can finish without intersection; let $y_{2}^{(R)}$ be the number of such jobs. None of the jobs with processing time 33 can finish without intersection.
@@
Consider the flow time by the large jobs. If we consider the jobs with processing time 22 or 33 and their intersections only, we have
@@
Similarly, consider the jobs with processing time 33 and their intersections only, we have
@@ 8.4 Deterministic Algorithms: Single Machine with Kill-and-Restart
   Release job 11 at 0 with processing time 44 ;
   Release job 22 at 22 with processing time 11 ;
 if ALG is working on job 11 before 33 then
      Release $c/2\cdot n$ jobs at 33 with processing time $\varepsilon$ ;
      Release $c/2\cdot n$ jobs at 77 with processing time $\varepsilon$ ;
@@
      Release $c\cdot n$ jobs at 55 with processing time $\varepsilon$ ;
@@
Against the adversary in Phase 11 of Algorithm 9, any deterministic online algorithm either incurs a flow time of $\Omega(n)$ or proceeds to Phase 2 with at least one of job 1 or job 2 unstarted. In all cases, OPT remains $O(1)$ and all jobs are solved.
@@
    If ALG is working on job 11 at time 33, then OPT could solve job 22 by 33 and job 11 by $7+c/2\cdot n\varepsilon$. For ALG, however:
@@
      If ALG keeps working on job 11 until finished, the flow time is at least $c/2\cdot n$;
@@
      If ALG kills at time 33 to solve $\varepsilon$-jobs, switch to job 22 and then switch to job 11:
@@
        If it doesn’t switch to $\varepsilon$-jobs at time 77, the flow time is at least $c/2\cdot n$;
@@
        Otherwise ALG must switch to $\varepsilon$-jobs at time 77, leaving job 11 unsolved.
@@
    Otherwise, ALG cannot finish job $1,2$ by time 55 while OPT can. In this case ALG would switch to job 22 at time 22 and then to solve job 11, or it would fall back to case 1.
@@
      If ALG keeps working on job 11 until finished, the flow time is at least $2c\cdot n$;
@@
      Otherwise ALG must switch to $\varepsilon$-jobs at time 55, leaving job 11 unsolved.
@@
Let N=n2cnN=n-2-cn be the total number of $\varepsilon$-jobs released in this phase. Consider any deterministic strategy employed by ALG. This strategy will result in some number of restarts of the large job $u$, say $r\geq 0$. These restarts partition the NN small jobs into $r+1$ groups. Let $a_{i}$ be the number of $\varepsilon$-jobs released during the $i$-th run of job $u$ (for $i=1,\dots,r+1$), where a run is a period of continuous work on $u$. We have $\sum_{i=1}^{r+1}a_{i}=N$.
@@ 8.5 Deterministic Algorithms: Multiple Machines without Kill-and-Restart
  • A job with processing time 11 is released at time 0.
@@
OPT can start all small jobs at their release time immediately; it can also solve the first job at 0 if $t>1$, or solve the first job at 22 otherwise. Hence, OPT is $O(1)$ in both cases. For ALG, as it can only use m1m-1 machines to solve the small jobs, at most m1m-1 jobs can be solved between the releases of two small job batchs. The total flow time for ALG is at least $\sum_{i=1}^{n/m}i/n=\Omega(n/m^{2})$. Therefore, the competitive ratio is $\Omega(n/m^{2})$.
@@ 9 Lower Bounds for Algorithms without the Knowledge of $n$
We now prove a contradiction: for any randomized algorithm ALG, there exists a constant $C>0$ such that for every positive integer $n_{0}$, one can construct an instance II with $|I|=n\geq n_{0}$ satisfying
@@
Adversary construction. We construct an oblivious adversary as follows. At time 0, the adversary releases a single job of size 22. For any algorithm, let $p(t)$ denote the probability density function that ALG starts this job at time $t$ if only the single job is present, which is known by the oblivious adversary since the probability distribution solely depend on the description of the algorithm and does not depend on algorithm’s randomized results. The algorithm must fall into one of the following categories (the parameters $x$ and $y$ will be fixed later):
@@
    In this case, the adversary releases $n_{1}$ dummy jobs of size 0 at time $t+1$. With probability at least $y$, ALG eventually starts the size-22 job within $[t,t+1)$, all these $n_{1}$ jobs complete no earlier than time $t+2$, so each contributes at least 11 to flow time and the algorithm’s total flow time is at least $n_{1}$, while an optimal solution completes all jobs by time 44, i.e., $\textup{{OPT}}\leq 4$. Therefore,
@@
    For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(t)\,dt<y$. In this case, the probability that the size-22 job has not been started by time $x$ is at least 1xy1-xy. The adversary releases $n_{0}$ dummy jobs of size 0 at time $x+1$. If ALG delays the size-22 job until after time $x$, each dummy job has flow time at least $x$, while $\textup{{OPT}}\leq 2$. Therefore,
@@
Let $x=n_{0}$ and $y=\frac{1}{2n_{0}}$, so that $xy=\frac{1}{2}$. Then, in the Type A case, there exists an instance II with $|I|=n\gg n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/(8n_{0}))\cdot n$; in the Type B case, there exists an instance II with $|I|=n=n_{0}+1>n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/8)\cdot n$.
@@ Appendix A Schedule Small Jobs: Non-Preemptive Shortest Job First
Define $t_{1}=A^{-1}(m(t+\tau))$ and let $t_{2}=t^{\prime}$. By Lemma 27, we have the volume bound: $V^{\prime}_{\leq p}(t_{1})\geq V_{\leq p}(t)$ for every pp.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t_{1}$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all ii.
@@
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 27. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all ii.

### v2

@@ 8.3 Generalization to Multiple Machines with Kill-and-Restart
In each busy gadget, if $c=0$, then schedule large job 0 at 0 and large job $1$ at $3$; if $c=1$, then schedule large job $1$ at 0 and large job 0 at $4$. In this way, no large jobs intersect with any small jobs. As a result, the total flow time in this gadget is at most $9$ (large jobs) +1+1 (fixed small jobs) +1+1 (random small jobs). Therefore, $\textup{{OPT}}\leq 11mk$ by summing up all gadgets. ∎
@@
Note that before time $2$, the algorithm does not know the realization of $c_{b}$ this round. Therefore, its behavior before time $2$ is not related to the result of $c_{b}$.
  • Let $x_{2}$ be the number of jobs with processing time $2$ started in [0.5, 2)[-0.5,\;2).
  • Let $x_{3}$ be the number of jobs with processing time $3$ started in [0.5, 2)[-0.5,\;2).
After time $2$, the algorithm scheduling decisions may depend on $c_{b}$. We consider the two cases separately:
  • Left case (w.p. $0.5$): None of the $x_{3}$ jobs can finish without intersection, and at most $x_{2}^{(L)}=\min\{m,x_{2}\}$ out of the $x_{2}$ jobs can finish without intersection. Let $y_{2}^{(L)}$ be the number of jobs with processing time $2$ starting in $[2,6.5)$ that finish without intersection, and $y_{3}^{(L)}$ be the number for processing time $3$. At most $m$ jobs out of $y_{2}^{(L)}+y_{3}^{(L)}$ can finish without intersection.
  • Right case (w.p. $0.5$): At most $m$ out of these $x_{2}+x_{3}$ jobs could end up without intersection with small jobs; let $x_{2}^{(R)}$ and $x_{3}^{(R)}$ be the number of such jobs, respectively. At most $m$ jobs with processing time $2$ released in $[2,6.5)$ can finish without intersection; let $y_{2}^{(R)}$ be the number of such jobs. None of the jobs with processing time $3$ can finish without intersection.
@@
Consider the flow time by the large jobs. If we consider the jobs with processing time $2$ or $3$ and their intersections only, we have
@@
Similarly, consider the jobs with processing time $3$ and their intersections only, we have
@@ 8.4 Deterministic Algorithms: Single Machine with Kill-and-Restart
   Release job $1$ at 0 with processing time $4$ ;
   Release job $2$ at $2$ with processing time $1$ ;
 if ALG is working on job $1$ before 33 then
      Release $c/2\cdot n$ jobs at $3$ with processing time $\varepsilon$ ;
      Release $c/2\cdot n$ jobs at $7$ with processing time $\varepsilon$ ;
@@
      Release $c\cdot n$ jobs at $5$ with processing time $\varepsilon$ ;
@@
Against the adversary in Phase $1$ of Algorithm 9, any deterministic online algorithm either incurs a flow time of $\Omega(n)$ or proceeds to Phase 2 with at least one of job 1 or job 2 unstarted. In all cases, OPT remains $O(1)$ and all jobs are solved.
@@
    If ALG is working on job $1$ at time $3$, then OPT could solve job $2$ by $3$ and job $1$ by $7+c/2\cdot n\varepsilon$. For ALG, however:
@@
      If ALG keeps working on job $1$ until finished, the flow time is at least $c/2\cdot n$;
@@
      If ALG kills at time $3$ to solve $\varepsilon$-jobs, switch to job $2$ and then switch to job $1$:
@@
        If it doesn’t switch to $\varepsilon$-jobs at time $7$, the flow time is at least $c/2\cdot n$;
@@
        Otherwise ALG must switch to $\varepsilon$-jobs at time $7$, leaving job $1$ unsolved.
@@
    Otherwise, ALG cannot finish job $1,2$ by time $5$ while OPT can. In this case ALG would switch to job $2$ at time $2$ and then to solve job $1$, or it would fall back to case 1.
@@
      If ALG keeps working on job $1$ until finished, the flow time is at least $2c\cdot n$;
@@
      Otherwise ALG must switch to $\varepsilon$-jobs at time $5$, leaving job $1$ unsolved.
@@
Let N=n2cnN=n-2-cn be the total number of $\varepsilon$-jobs released in this phase. Consider any deterministic strategy employed by ALG. This strategy will result in some number of restarts of the large job $u$, say $r\geq 0$. These restarts partition the $N$ small jobs into $r+1$ groups. Let $a_{i}$ be the number of $\varepsilon$-jobs released during the $i$-th run of job $u$ (for $i=1,\dots,r+1$), where a run is a period of continuous work on $u$. We have $\sum_{i=1}^{r+1}a_{i}=N$.
@@ 8.5 Deterministic Algorithms: Multiple Machines without Kill-and-Restart
  • A job with processing time $1$ is released at time 0.
@@
OPT can start all small jobs at their release time immediately; it can also solve the first job at 0 if $t>1$, or solve the first job at $2$ otherwise. Hence, OPT is $O(1)$ in both cases. For ALG, as it can only use m1m-1 machines to solve the small jobs, at most m1m-1 jobs can be solved between the releases of two small job batchs. The total flow time for ALG is at least $\sum_{i=1}^{n/m}i/n=\Omega(n/m^{2})$. Therefore, the competitive ratio is $\Omega(n/m^{2})$.
@@ 9 Lower Bounds for Algorithms without the Knowledge of $n$
We now prove a contradiction: for any randomized algorithm ALG, there exists a constant $C>0$ such that for every positive integer $n_{0}$, one can construct an instance $I$ with $|I|=n\geq n_{0}$ satisfying
@@
Adversary construction. We construct an oblivious adversary as follows. At time 0, the adversary releases a single job of size $2$. For any algorithm, let $p(t)$ denote the probability density function that ALG starts this job at time $t$ if only the single job is present, which is known by the oblivious adversary since the probability distribution solely depend on the description of the algorithm and does not depend on algorithm’s randomized results. The algorithm must fall into one of the following categories (the parameters $x$ and $y$ will be fixed later):
@@
    In this case, the adversary releases $n_{1}$ dummy jobs of size 0 at time $t+1$. With probability at least $y$, ALG eventually starts the size-$2$ job within $[t,t+1)$, all these $n_{1}$ jobs complete no earlier than time $t+2$, so each contributes at least $1$ to flow time and the algorithm’s total flow time is at least $n_{1}$, while an optimal solution completes all jobs by time $4$, i.e., $\textup{{OPT}}\leq 4$. Therefore,
@@
    For all $t\in\{1,2,\dots,x\}$, it holds that $\int_{t}^{t+1}p(t)\,dt<y$. In this case, the probability that the size-$2$ job has not been started by time $x$ is at least 1xy1-xy. The adversary releases $n_{0}$ dummy jobs of size 0 at time $x+1$. If ALG delays the size-$2$ job until after time $x$, each dummy job has flow time at least $x$, while $\textup{{OPT}}\leq 2$. Therefore,
@@
Let $x=n_{0}$ and $y=\frac{1}{2n_{0}}$, so that $xy=\frac{1}{2}$. Then, in the Type A case, there exists an instance $I$ with $|I|=n\gg n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/(8n_{0}))\cdot n$; in the Type B case, there exists an instance $I$ with $|I|=n=n_{0}+1>n_{0}$ such that $\mathbb{E}[\textup{{ALG}}]/\textup{{OPT}}>(1/8)\cdot n$.
@@ Appendix A Schedule Small Jobs: Non-Preemptive Shortest Job First
Define $t_{1}=A^{-1}(m(t+\tau))$ and let $t_{2}=t^{\prime}$. By Lemma 27, we have the volume bound: $V^{\prime}_{\leq p}(t_{1})\geq V_{\leq p}(t)$ for every $p$.
We claim the following: the $i$-th smallest job (in terms of size) that has started (either in processing or completed) under NSJF by time $t_{1}$, denoted by size $p^{\prime}_{i}$, must not exceed the size of the $i$-th smallest completed job under OPT by time $t$, denoted $p_{i}$. That is, we aim to prove: $p^{\prime}_{i}\leq p_{i}$ for all $i$.
@@
where the second inequality uses the assumption that $p^{\prime}_{j}\leq p_{j}$ for all $j<i$. Combining these two, we get: $V^{\prime}_{\leq p_{i}}(t_{1})<V_{\leq p_{i}}(t),$ which contradicts the earlier volume bound from Lemma 27. Hence, our assumption must be false, and we conclude: $p^{\prime}_{i}\leq p_{i}$ for all $i$.